An electrical circuit has three devices each plugged into a standard household powerpoint 230V. Device A has a resistance of 1.0 kΩ (kiloohm), device B has a resistance of 3.0 kΩ and device C has a resistance of 6.0 kΩ.
If all three devices are turned on for 15 minutes, what will the total energy consumption for all three devices.
1
Expert's answer
2020-05-05T18:39:52-0400

Total resistance of all devices

"R=\frac{R_1\times R_2+R_2\times R_3+R_1\times R_3}{R_1\times R_2\times R_3};\\R=\frac{1\times 3+3\times 6+1\times 6}{1\times 3\times 6}=1.5k\Omega=1500\Omega;15min=15\times60=900s"

The total energy consumption

"A=\frac{U^2}{R}\times t=\frac{230^2}{1500}\times 900=31740J=31.740kJ"

​ Answer: The total energy consumption is 31.740kJ